# Exploring the Design of GenAI-Based Systems to Support Socially Shared Metacognition

### 09 Mar 2026 v1

@@ 1. Introduction
Yet SSM rarely emerges spontaneously (Tsai et al., 2018; Volet et al., 2009a), creating needs for intervention (Silva et al., 2023; Panadero and Järvelä, 2015). Thanks to their understanding and generative capacity (Zhao et al., 2026; Hu et al., 2025b; Zhao, 2025; Zhao et al., ; Zhao et al., 2024b, a), GenAI-based systems present new possibilities for supporting SSM through, for example, automated monitoring of group processes and timely feedback provision (George and Wooden, 2023; Song and Xiong, 2021). However, poorly designed GenAI-based systems may encourage over-reliance on AI-generated explicit instruction and erode groups’ capacity to develop autonomous regulatory processes (Liu et al., 2023; Yang et al., 2025; Kim et al., 2025; Bauer et al., 2025).

### 11 Mar 2026 v2

@@ 1. Introduction
Yet SSM rarely emerges spontaneously (Tsai et al., 2018; Volet et al., 2009a), creating needs for intervention (Silva et al., 2023; Panadero and Järvelä, 2015). Thanks to their understanding and generative capacity (Zhao et al., 2026a; Hu et al., 2025b; Zhao, 2025; Zhao et al., ; Zhao et al., 2024c, b), GenAI-based systems present new possibilities for supporting SSM through, for example, automated monitoring of group processes and timely feedback provision (George and Wooden, 2023; Song and Xiong, 2021). However, poorly designed GenAI-based systems may encourage over-reliance on AI-generated explicit instruction and erode groups’ capacity to develop autonomous regulatory processes (Liu et al., 2023; Yang et al., 2025; Kim et al., 2025; Bauer et al., 2025).
@@ 5. Discussion
We hope to discuss these considerations in this workshop, and the expertise we can potentially bring spans machine learning, AI-based systems for individual and collaborative knowledge and creative work, interaction design with AI, and cognitive and metacognitive augmentation (Zhao et al., 2026a; Zhao, 2025; Zhao et al., ; Zhao et al., 2024c, b, 2026b, a; Hu et al., 2025b; Lisena et al., 2026; Simperl et al., 2025a, b; Zhang et al., 2025; Zhao, 2021; Luo et al., 2021).